OpenStudy (anonymous):

help im so bad at math its not even funny! Explain the difference between x + x and x∙x. How will the answer differ and how do they simplify differently?

OpenStudy (anonymous):

What is 2 + 2? What is 2 * 2?

OpenStudy (anonymous):

2+2= 4 2*2= 4

OpenStudy (anonymous):

Lol bad example. Whats 3+3? Whats 3*3?

OpenStudy (anonymous):

hahah 3+3=6 3*3=9

OpenStudy (anonymous):

\[3*3 = 3^2 = 9\] right?

OpenStudy (anonymous):

very true

OpenStudy (anonymous):

So then\[x*x = x^2\]

OpenStudy (anonymous):

ohh okay i get it now!

OpenStudy (anonymous):

Thats good! This is how it simplifies differently, you can use the 3 + 3 and 3*3 example for how they answer differently :)

OpenStudy (anonymous):

thank you for helping me so much! haha :)

OpenStudy (anonymous):

Yw